Question 9

Which of the following is not a step in hypothesis testing?
(1) Formulate the null and alternative hypothesis.
(2) Specify the level of significance.
(3) Determine the test statistic, critical value and p-value.
(4) Determine the rejection region.
(5) None of the above.

Question 11

What is the value of the test statistic?
(1) -3.33
(2) 2.58
(3) 1.96
(4) 1.645
(5) 0.0012

Question 12

What is the critical value?
(1) -3.33
(2) 2.58
(3) 1.96
(4) 1.645
(5) 0.0012

Question 13

What is the p-value?
(1) 0.05
(2) 0.01
(3) 0.0024
(4) p–value < 0.002
(5) None of the above
1
Expert's answer
2020-03-17T12:26:27-0400

Question 9 

Which of the following is not a step in hypothesis testing?

(1) Formulate the null and alternative hypothesis.

(2) Specify the level of significance.

(3) Determine the test statistic, critical value and p-value.

(4) Determine the rejection region.

(5) None of the above.

Solution

(5) None of the above.


Botlhale, a practicing Statistician is tasked with investigating whether the population mean of a random variable "X" is different from "100." She has the following additional information to help her with the investigation: "\alpha=0.01, \sigma=3, \bar {X}=99," and "n=100."

Question 11

What is the value of the test statistic?

(1) -3.33

(2) 2.58

(3) 1.96

(4) 1.645

(5) 0.0012 

Solution


"Z={\bar{X}-\mu \over \sigma/\sqrt{n}}={99-100\over 3 \sqrt{100}}\approx-3.33"

(1) -3.33


Question 12

What is the critical value?

(1) -3.33

(2) 2.58

(3) 1.96

(4) 1.645

(5) 0.0012 

Solution

The z-critical values for a two-tailed test, for a significance level of "\alpha=0.01:z_c=-2.58" and "z_c=2.58"

(2) 2.58


Question 13

What is the p-value?

(1) 0.05

(2) 0.01

(3) 0.0024

(4) p–value < 0.002

(5) None of the above

Solution

"p-"Value from Z Score Calculator: "Z=-3.33, \alpha=0.01," two-tailed

The "p" -value is "0.000868"

(4) p–value < 0.002